OpenStudy (anonymous):

how do I deal with the square power in this problem: Integrate: [ln(x))^2] w/ U-sub or IBP Am I able to use U-Sub if u=ln(x) ?

OpenStudy (waynex):

IBP is my first thought, since it's the analog of the product rule. If you use u=ln(x), what is du? It's 1/x dx, and 1/x is not in that integral. u substitution doesn't look promising to me.

OpenStudy (anonymous):

It's almost like this -- please look at the picture below

OpenStudy (waynex):

alexuspanait, that rule for logarithms applies to \[\ln(x^a) = a \cdot \ln(x).\] Note the difference between the following: \[\ln(x^a) \neq \ln(x)^a.\]

OpenStudy (anonymous):

@Waynex - you're right , I was wrong , sorry, my bad I tried to help ...

OpenStudy (waynex):

There is no need to be sorry. I only mentioned it to help avoid any confusion. Don't let that deter you. Best

OpenStudy (anonymous):

Well if you know you log properties which is the same for natural logs, then \[\ln a ^b = b \ln a\] so you end up getting... \[\int\limits_{?}^{?} \ln x ^2 = \int\limits_{?}^{?} 2 \ln x\]

OpenStudy (anonymous):